# Self-bound quark stars with a first-order two-to-three flavor phase transition

### v1

@@ VI.3 Gravitational versus baryonic compactness
To quantify this behavior, Fig. 12(a) includes the analytic fit introduced in Ref. [Lugones2025universality], which accurately describes the combined vMIT and flavor-blind QMDD strange-star dataset. The accuracy is assessed in Fig. 12(b), which presents the relative deviation from that fit. For many parameterizations the scatter remains below the percent level, whereas for others it can reach up to $\sim 20\%$, mirroring the residual patterns seen in the $\bar{I}$$\mathscr{C}$ and $\Lambda$$\mathscr{C}$ cases.
@@ VII Conclusions
    Phase structure and stability. For broad regions of $(a,C)$, the flavor–dependent EOS exhibits a first–order $ud\!\to\!uds$ phase transition whose transition pressure decreases with increasing excluded–volume parameter $\kappa$. The associated energy–density jumps satisfy Seidov’s bound below the instability threshold. Quantitatively, the ratio $(\epsilon_{+}/\epsilon_{-})\big/\bigl[\tfrac{3}{2}(1+p_{\mathrm{tr}}/\epsilon_{-})\bigr]$ lies in the range 0.850.850.970.97 for all cases considered and is essentially insensitive to $\kappa$ at fixed $(a,C)$ (cf. Eq. (6) and Table 2). In line with this expectation, the hybrid self-bound configurations occupy stable branches of the mass–radius relation.
@@
    Global stellar properties and observational constraints. The parameter $\kappa$ has a strong and systematic impact on the stiffness of the EOS and therefore on masses, radii, tidal deformabilities, and moments of inertia. In the absence of excluded volume ($\kappa=0$), maximum masses tend to undershoot the $\sim 2\,M_{\odot}$ constraint, whereas large $\kappa$ values can overshoot radius bounds at canonical masses. Intermediate repulsion (e.g., $\kappa\!\approx\!0.3$) often reconciles $M_{\max}\!\gtrsim\!2\,M_{\odot}$ with recent NICER inferences [Miller2019, Riley2019, Miller2021, Riley2021] and the compact object in HESS J1731-347 [Doroshenko2022]. Consistent with our detailed results, $\Lambda(M)$ decreases steeply with mass and increases with $\kappa$; self–bound hybrid stars are typically more compatible with the GW170817 constraint at $1.4\,M_{\odot}$ than pure SQM stars, the latter requiring comparatively strong repulsion to approach the observed range.
@@
    Universal relations. Three EOS–insensitive trends emerge: (a) the dimensionless moment of inertia $\bar{I}\!\equiv\!I/M_{G}^{3}$ versus compactness $\mathscr{C}$, (b) the tidal deformability $\Lambda$ versus $\mathscr{C}$, and (c) the gravitational versus baryonic compactness, $\mathscr{C}$$\mathscr{C}_{B}$. In all cases, the explicit dependence on $\kappa$ largely disappears once the relations are expressed in dimensionless form. These relations follow the flavor–blind strange–quark–star fits reported previously [Lugones2025universality], with typical residuals below the percent level and larger deviations up to $\sim 10$$20\%$ for a few $a$ families (where the phase transition reshapes the stellar structure). The $\mathscr{C}$$\mathscr{C}_{B}$ mapping is less tight. Self-bound hybrid stars remain close to the universal relations but exhibit slightly larger departures than strange quark stars, due to the change in stiffness across the two- to three-flavor EOS transition. Because the onset of this transition varies among parameter sets, the effect manifests as a spread of curves for some parametrizations.
@@
    Quark–hadron discrimination. For fixed compactness, self–bound stars generically yield smaller Love numbers and smaller $\Lambda$ than hadronic stars, and their $\bar{I}$$\mathscr{C}$ curves also differ from hadronic fits. Hence, joint constraints on $(M_{G},R)$ together with either $I$ or $\Lambda$ can help discriminate self–bound quark stars from purely hadronic stars. We also find that models with $\kappa\!=\!0$ yield $I(M)$ closer to hadronic benchmarks (SLy, APR), whereas increasing $\kappa$ systematically raises II—a trend that future moment–of–inertia measurements could exploit.
Overall, the flavor–dependent QMDD model with excluded volume accommodates both SQM stars and self–bound hybrid stars compatible with current multimessenger constraints, and it exhibits EOS–insensitive trends that are useful for inference. The relations involving $\bar{I}$ and $\Lambda$ are comparatively tight, whereas the $\mathscr{C}$$\mathscr{C}_{B}$ mapping, though practical for first–order estimates of baryonic masses and binding energies, carries larger model–dependent scatter in specific regions of parameter space.
@@ A.2 Excluded–volume effects
We incorporate short–range repulsion via an excluded–volume prescription in the $T=0$ Helmholtz representation, following Refs. [Lugones:2023zfd, Lugones:2024ryz]. Consistent with the flavor–dependent mass model, we take
@@ B.2 Dimensionless tidal deformability
where $c_{s}^{2}\equiv dp/d\epsilon$, and $m(r)$, $p(r)$, $\epsilon(r)$ are the TOV solutions from Eqs. (28)–(30). Regularity at the center sets $y(0)=2$. The Love number $k_{2}$ and hence $\Lambda$ follow by integrating (38) together with the TOV equations up to $r=R$. For additional details, see Ref. [Postnikov:2010yn].
Junction conditions at a first–order interface.
When a sharp ($ud\!\to\!uds$) first–order transition occurs at pressure $p_{\rm tr}$, located at radius $r_{\rm tr}$ defined by $p(r_{\rm tr})=p_{\rm tr}$, the stellar solution must satisfy the Israel–Darmois junction conditions across the infinitesimal interface (assumed free of surface stresses). In static spherical symmetry these reduce to continuity of the induced metric and extrinsic curvature, which in practice implies continuity of the metric functions (hence of $m$ and $\nu$) and of the pressure, while allowing a jump of the energy density from $\epsilon_{-}$ to $\epsilon_{+}$ (with $\Delta\epsilon\equiv\epsilon_{+}-\epsilon_{-}>0$). For the tidal sector, the first–order ODE for $y(r)$ must be matched by enforcing the standard jump obtained by integrating the perturbation equation across the density discontinuity ($\delta\rightarrow 0$) [Postnikov:2010yn, Takatsy:2020bnx]:
$y(r_{\rm tr}+\delta)=y(r_{\rm tr}-\delta)-\frac{4\pi\,r_{\rm tr}^{3}\,\Delta\epsilon}{\,m(r_{\rm tr})+4\pi r_{\rm tr}^{3}p_{\rm tr}\,}\,$, (41)
after which the integration proceeds to the surface to compute $k_{2}$ and $\Lambda$. This prescription guarantees continuity of the physical (gauge–invariant) metric perturbations across the interface and a well–posed Love number.
@@ B.3 Moment of inertia for slow rigid rotation
We evaluate the stellar moment of inertia within the slow–rotation (Hartle–Thorne) framework, restricting attention to uniform rotation with angular frequency $\Omega$ as measured at infinity [Hartle:1967he]. To first order in $\Omega$, the total angular momentum is $J=\Omega\,I$, so that
@@
with $R$ the radius where $p(R)=0$. This formulation is equivalent to the standard computation based on the frame–dragging function $\bar{\omega}(r)=\Omega-\omega(r)$ and its exterior matching; see, e.g., Refs. [Hu:2023vsq, Dong:2023vxv].

### v2

@@ VI.2 Gravitational versus baryonic compactness
To quantify this behavior, Fig. 11(a) includes the analytic fit introduced in Ref. [19], which accurately describes the combined vMIT and flavor-blind QMDD strange-star dataset. The accuracy is assessed in Fig. 11(b), which presents the relative deviation from that fit. For many parameterizations the scatter remains below the percent level, whereas for others it can reach up to $\sim 20\%$, mirroring the residual patterns seen in the $\bar{I}$$\mathscr{C}$ and $\Lambda$$\mathscr{C}$ cases.
@@ VII Conclusions
    Phase structure and stability. Over broad regions of the $(a,C)$ plane, the flavor–dependent QMDD EOS develops a first–order $ud\!\to\!uds$ transition. For the benchmark families analyzed in detail, the transition pressure $p_{\rm tr}$ shifts to lower values as the excluded–volume parameter $\kappa$ is increased. The corresponding energy–density discontinuities remain below Seidov’s onset–instability threshold: the Seidov diagnostic $\gamma\equiv(\epsilon_{+}/\epsilon_{-})\big/\bigl[\tfrac{3}{2}\bigl(1+p_{\rm tr}/\epsilon_{-}\bigr)\bigr]$ lies in the range 0.850.850.970.97 for all cases reported in Table 2 (cf. Eq. (6)), and is essentially insensitive to $\kappa$ at fixed $(a,C)$. Consistently, the full general–relativistic equilibrium sequences exhibit stable hybrid self–bound branches (identified by the standard turning–point criterion along each one–parameter family) in the mass–radius relation.
@@
    Global stellar properties and observational constraints. Over a broad region of the EOS parameter space, the two-to-three flavor transition sets in before the maximum-mass configuration is reached, i.e., $p_{c}^{\max}\!\geq\!p_{\rm tr}$ (equivalently, the udsuds core forms along the stable branch prior to $M_{\max}$). As a result, self-bound hybrid stars constitute a generic outcome of the flavor-dependent scheme, and the onset of the udsuds core produces a characteristic kink in the stellar sequences at $p_{c}=p_{\rm tr}$, visible in the MMRR, $\Lambda(M)$, and $I(M)$ curves. The parameter $\kappa$ has a strong and systematic impact on the stiffness of the EOS and therefore on masses, radii, tidal deformabilities, and moments of inertia. In the absence of excluded volume ($\kappa=0$), maximum masses tend to undershoot the $\sim 2\,M_{\odot}$ constraint, whereas large $\kappa$ values can overshoot radius bounds at canonical masses. The tidal deformability provides an additional and complementary filter. Comparing our $\Lambda(M)$ curves with the GW170817 low-spin upper limit $\Lambda(1.4\,M_{\odot})\leq 800$ [1], we find that pure strange-quark-star sequences satisfy the bound for all explored values of $\kappa$, owing to their intrinsically high compactness. For self-bound hybrid stars the constraint is more selective: at low and intermediate transition densities, only the $\kappa=0$ models comply with the tidal bound, and they do so only marginally—yet these same parametrizations fail to reach $2\,M_{\odot}$, effectively ruling out those EOS sets. The tension is resolved for high-transition-density configurations, where $\kappa=0.3$ simultaneously satisfies $\Lambda(1.4\,M_{\odot})\lesssim 800$ and $M_{\max}\gtrsim 2\,M_{\odot}$, while also remaining consistent with NICER radius inferences [23, 28, 24, 29] and the compact object in HESS J1731-347 [6]. Taken together, the mass, radius, and tidal constraints converge on intermediate repulsion ($\kappa\approx 0.3$) with a high transition density as the most observationally viable region of parameter space for self-bound hybrid configurations.
@@
    Universal relations. Two EOS–insensitive trends emerge: (a) the dimensionless moment of inertia $\bar{I}\!\equiv\!I/M_{G}^{3}$ versus compactness $\mathscr{C}$, and (b) the gravitational versus baryonic compactness, $\mathscr{C}$$\mathscr{C}_{B}$. In both cases, the explicit dependence on $\kappa$ largely disappears once the relations are expressed in dimensionless form. These relations follow the flavor–blind strange–quark–star fits reported previously [19], with typical residuals below the percent level and larger deviations up to $\sim 10$$20\%$ for a few $a$ families (where the phase transition reshapes the stellar structure). The $\mathscr{C}$$\mathscr{C}_{B}$ mapping is less tight. Self-bound hybrid stars remain close to the universal relations but exhibit slightly larger departures than strange quark stars, due to the change in stiffness across the two- to three-flavor EOS transition. Because the onset of this transition varies among parameter sets, the effect manifests as a spread of curves for some parametrizations.
@@
    Quark–hadron discrimination. For fixed compactness, self–bound stars follow $\bar{I}$$\mathscr{C}$ systematics that can differ from hadronic fits, reflecting their distinct surface boundary condition (finite surface density) and more homogeneous interiors. In practice, however, the separation is not uniform: in the high–compactness regime most relevant for heavy stars the differences with respect to hadronic trends can become comparatively modest. Anyway, joint constraints on $(M,R)$ together with an independent moment–of–inertia measurement may help discriminate self–bound quark stars from purely hadronic stars, but this test will likely require sufficiently precise data. We also find that models with $\kappa=0$ yield $I(M)$ closer to representative hadronic benchmarks (e.g., SLy and APR), whereas increasing $\kappa$ systematically raises $I$; this trend could be leveraged by future moment–of–inertia measurements.
Overall, the flavor–dependent QMDD model with excluded volume accommodates both SQM stars and self–bound hybrid stars compatible with current astrophysical constraints, and it exhibits EOS–insensitive trends that can be useful for inference within the self–bound sector. In particular, the $\bar{I}$$\mathscr{C}$ relation is comparatively tight across our parameter space, whereas the $\mathscr{C}$$\mathscr{C}_{B}$ mapping—though practical for first–order estimates of baryonic masses and binding energies—shows a larger, model–dependent scatter in specific regions of parameter space.
@@ A.2 Excluded–volume effects
We incorporate short–range repulsion via an excluded–volume prescription in the $T=0$ Helmholtz representation, following Refs. [18, 20]. Consistent with the flavor–dependent mass model, we take
@@ B.2 Dimensionless tidal deformability
where $c_{s}^{2}\equiv dp/d\epsilon$, and $m(r)$, $p(r)$, $\epsilon(r)$ are the TOV solutions from Eqs. (28)–(30). Regularity at the center sets $y(0)=2$. The Love number $k_{2}$ and hence $\Lambda$ follow by integrating (38) together with the TOV equations up to $r=R$. For additional details, see Ref. [27].
Junction conditions at a sharp interface.
When a sharp ($ud\!\to\!uds$) first–order transition occurs at pressure $p_{\rm tr}$, the interface is located at the radius $r_{\rm tr}$ defined by $p(r_{\rm tr})=p_{\rm tr}$. Assuming an infinitesimally thin interface free of surface stresses, the Israel–Darmois junction conditions imply continuity of the induced metric and extrinsic curvature. In static spherical symmetry this translates into continuity of the metric functions (hence $m$ and $\nu$) and of the pressure $p$, while allowing a jump in the energy density.
To avoid any sign ambiguity, we define the density jump as
$\epsilon_{\rm in}\equiv\epsilon(r_{\rm tr}-0),\quad\epsilon_{\rm out}\equiv\epsilon(r_{\rm tr}+0),\quad\Delta\epsilon\equiv\epsilon_{\rm in}-\epsilon_{\rm out}>0,$ (41)
where “in” (“out”) denotes the inner (outer) side of the interface, i.e. smaller (larger) $r$. For the tidal sector, the first–order ordinary differential equation for $y(r)=rH^{\prime}(r)/H(r)$ must be matched by integrating the perturbation equation across the density discontinuity [27, 31], which yields
$y(r_{\rm tr}+0)=y(r_{\rm tr}-0)-\frac{4\pi\,r_{\rm tr}^{3}\,\Delta\epsilon}{\,m(r_{\rm tr})+4\pi r_{\rm tr}^{3}p_{\rm tr}\,}\,.$ (42)
After enforcing (42), the integration proceeds to the stellar surface in order to compute $k_{2}$ and $\Lambda$. This prescription ensures continuity of the physical (gauge–invariant) metric perturbations and a well–posed Love number.
Surface correction for self–bound stars.
For self–bound configurations with a finite surface density $\epsilon_{s}\equiv\epsilon(R^{-})>0$, the pressure vanishes at the surface, $p(R)=0$, while the exterior is vacuum, $\epsilon(R^{+})=0$. Equation (42) then applies with $r_{\rm tr}=R$, $p_{\rm tr}=0$, $m(R)=M$, and $\Delta\epsilon=\epsilon_{s}$, giving
$y(R^{+})=y(R^{-})-\frac{4\pi R^{3}\,\epsilon_{s}}{M}\,.$ (43)
The value entering the standard $k_{2}(\beta,y_{R})$ formula is the exterior one, $y_{R}\equiv y(R^{+})$.
@@ B.3 Moment of inertia for slow rigid rotation
We evaluate the stellar moment of inertia within the slow–rotation (Hartle–Thorne) framework, restricting attention to uniform rotation with angular frequency $\Omega$ as measured at infinity [8]. To first order in $\Omega$, the total angular momentum is $J=\Omega\,I$, so that
@@
with $R$ the radius where $p(R)=0$. This formulation is equivalent to the standard computation based on the frame–dragging function $\bar{\omega}(r)=\Omega-\omega(r)$ and its exterior matching; see, e.g., Refs. [10, 5].